# Stable generative modeling using diffusion maps

### v1

@@ 3. Diffusion approximation
(8) $p_{\epsilon}(x)=\frac{D(v_{\epsilon})t_{\epsilon}(x)}{v_{\epsilon}^{\rm T}t_{\epsilon}(x)}$
for all $x\in\mathbb{R}^{d}$. This vector gives the transition probabilities from $x$ to $\{x^{(i)}\}_{i=1}^{M}$ and provides a finite-dimensional approximation to the conditional probability distribution $\pi_{\epsilon}(\cdot|x)$ of the true underlying diffusion process; that is, the semigroup $\exp(\epsilon\mathcal{L})$ with generator $\mathcal{L}$. See Section 4 for more details. In the following, we further approximate $\pi_{\epsilon}(\cdot|x)$ by a Gaussian and estimate its mean and covariance matrix using the probability vector $p_{\epsilon}$ and the data matrix of samples
(9) $\mathcal{X}=(x^{(1)},\ldots,x^{(M)})\in\mathbb{R}^{d\times M}.$
3.1. Sampling algorithms
We now present our main Langevin sampling strategies based on the diffusion map approximation for the probability vector $p_{\epsilon}(x)$ in (8). Our sampling schemes have the same drift term, $m_{\epsilon}(x)$, but differ in the way the diffusion matrix, $\Sigma(x)$, is defined. We consider a data-unaware diffusion as well as a data-aware diffusion which turns out to be advantageous in generating new samples from the data distribution $\pi$. See the numerical experiments in Section 6.
3.1.1. Langevin sampler with data-unaware diffusion
Following the proposed methodology, we introduce the sample-based approximation of the conditional mean
@@ 3.1.1. Langevin sampler with data-unaware diffusion
(11) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m_{\epsilon}(X_{n})-X_{n}}{\epsilon}\right)+\sqrt{2\Delta\tau}K(X_{n})^{1/2}\Xi_{n},$
where $\Delta\tau$ is the time step. In other words, we obtain the score function approximation
@@
in (1). By taking $\Delta\tau=\epsilon$, we have
(13) $X_{n+1}=m_{\epsilon}(X_{n})+\sqrt{2\epsilon}K(X_{n})^{1/2}\Xi_{n}.$
Note that (13) fits into the general formulation (3) with $\Sigma(x)=\sqrt{2\epsilon}K(x)^{1/2}$.
The parameter $\epsilon>0$ can be seen as a step-size. For large $\epsilon$, the expected value $m_{\epsilon}(x)$ will become essentially independent of the current state $X_{n}$ and the diffusion process will sample from a centred Gaussian. For $\epsilon\to 0$, on the other hand, the probability vector $p_{\epsilon}(x)$ can potentially degenerate into a vector with a single entry approaching one with all other entries essentially becoming zero. Hence a key algorithmic challenge is to find a good value for $\epsilon$ and a suitable $K(x)$, which guarantee both good mixing and accuracy, that is, $X_{n}\sim\pi$ as $n\to\infty$.
@@ 3.1.2. Langevin sampler with data-aware diffusion
From (8) and (9), one can also define the conditional covariance matrix,
(14) $C(x)=(\mathcal{X}-m_{\epsilon}(x)1_{\rm M}^{\rm T})D(p_{\epsilon}(x))(\mathcal{X}-m_{\epsilon}(x)1_{\rm M}^{\rm T})^{\rm T}\in\mathbb{R}^{d\times d},$
which is the covariance matrix associated with the probability vector $p_{\epsilon}(x)$. Therefore, one can more directly implement a Gaussian approximation associated with the transition probabilities $p_{\epsilon}(x)$ and introduce the update
(15) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m_{\epsilon}(X_{n})-X_{n}}{\epsilon}\right)+\sqrt{\Delta\tau/\epsilon}C(X_{n})^{1/2}\Xi_{n}.$
@@
(16) $X_{n+1}=m_{\epsilon}(X_{n})+C(X_{n})^{1/2}\Xi_{n},$
which we found to work rather well in our numerical experiments since it directly captures the uncertainty contained in the data-driven coupling $P_{\epsilon}$. The scheme (16) corresponds to setting $\Sigma(x)=C(x)^{1/2}$ in (3). Also note that the schemes (16) still depends on $K(X)$ through the probability vector $p_{\epsilon}(x)$.
It is not always justifiable to use $\Delta\tau=\epsilon$ as a step-size and $\Delta\tau<\epsilon$ can be beneficial instead. In those cases, one can resort to the re-scaled time-stepping methods (11) or (15), respectively.
@@ 3.2. Algorithmic properties
Follows from the definition (10) and the fact that $p_{\epsilon}(x)$ is a probability vector for all $\epsilon>0$ and all $x\in\mathbb{R}^{d}$. ∎
This establishes stability of the Langevin samplers (13) and (16) for all step-sizes $\epsilon>0$.
The next lemma shows that the Langevin sampler (13) is geometrically ergodic toward the invariant measure $\pi$.
@@
$C=\{x:\|x\|\leq R\}$
for suitable $R>0$. Since $m_{\epsilon}(X_{n})\in\mathcal{C}_{M}$ and $\pi$ has compact support, one can find a radius $R>0$, which is independent of the training data $\{x^{(i)}\}$, such that $\mathcal{C}_{M}\subset C$ and
@@
for all $X_{n}\notin C$ with $10\leq\lambda<$ 1. Furthermore, because of the additive Gaussian noise in (13), there is a probability density function $\nu(x)$ and a constant $\delta>0$ such that
@@
for all $x,x^{\prime}\in C$. Here ${\rm n}(x;m,\Sigma)$ denotes the Gaussian probability density function with mean $m$ and covariance matrix $\Sigma$. In other words, $C$ is a small set in the sense of [35]. Geometric ergodicity follows from Theorem 15.0.1 in [35]. See also the self-contained presentation in [33]. ∎
Extending Lemma 3.2 to the time-stepping scheme (16) is non-trivial since the covariance matrix (14) may become singular.
Lemma 3.1 also suggests to replace the sampling step (13) by the associated split-step scheme
(18a) $X_{n+1/2}$ $=X_{n}+\sqrt{2\epsilon}K(X_{n})^{1/2}\Xi_{n},$
(18b) $X_{n+1}$ $=m_{\epsilon}(X_{n+1/2}).$
@@ 3.3. Variable bandwidth diffusion
It is well-known from the literature on diffusion maps that a variable bandwidth can improve the approximation quality for fixed sample size $M$ [1]. Here we utilize the same idea. However, we no longer insist on approximating the standard generator with $K=I$, since we only wish to sample from the distribution $\pi$ rapidly. Hence, we consider (4) with $K(x)$ of the form
@@
It is an active area of research to select a $\rho$ that increases the spectral gap of $\mathcal{L}$ while not increasing computational complexity. Indeed, a larger spectral gap implies a faster convergence rate [41], indicating that the generated samples are closer to the reference at a finite time, exhibiting a high accuracy. We demonstrate numerically in Section 6 that $\rho$ can indeed be used to increase the sampling accuracy. More specifically, the bandwidth $\rho(x)$ is chosen as
@@
where $\beta\leq 0$ is a parameter and the unknown sampling distribution $\pi$ is approximated by an inexpensive low accuracy density estimator. One finds that the variable bandwidth parameter $\beta$ in (21) and the scaling parameter $\epsilon$ both influence the effective step-size in the reference process  (4) for (20). In order to disentangle the two scaling effects we modify the construction of the entries (5) in $T_{\epsilon}$ as follows. We first compute $\pi_{i}\approx\pi(x^{(i)})$ over all data points and then rescale these typically unnormalized densities:
@@
for $i=1,\ldots,M$, that is, $K(x^{(i)})=\rho_{i}I$ in (5) and, more generally,
@@ 4. Connections to the generator of Langevin dynamics
In this section, we discuss the proposed methodology in the limit $M\to\infty$ under the idealised assumption that the target distribution $\pi$ is in fact known. It is well-known that $K=I$ implies that the diffusion map approximates the semi-group corresponding to the generator of standard Langevin dynamics. Employing a variable bandwidth (20) instead, i.e. $K=\rho(x)I$, alters the semi-group and thus the underlying generator. However, since in both cases the density $\pi(x)$ remains invariant, this is not an issue as we are only concerned with drawing samples from the distribution and not in reproducing any dynamical features.
More precisely, using (20) in (4), we consider the diffusion map approximation given by (6). We formally take the limit $M\to\infty$ and denote the limiting operator by $\mathcal{P}_{\epsilon}$ [58]. One formally obtains
@@
(23) $\mathcal{L}f=\pi^{-1}\nabla\cdot(\pi\rho\nabla f)=\nabla\cdot(\rho\nabla f)+\rho\nabla\log\pi\cdot\nabla f$
@@
$\dot{X}_{t}=\rho(X_{t})\left(\nabla(\log\pi(X_{t})-\log\mu(X_{t})\right),\qquad X_{0}\sim\mu_{0},$
@@
Please also note that the generator corresponds to the diffusion process
(24) $\dot{X}_{t}=\rho(X_{t})\nabla\log\pi(X_{t})+\nabla\rho(X_{t})+\sqrt{2\rho(X_{t})}\dot{W}$
in Itô form and to
$\dot{X}_{t}=\rho(X_{t})\nabla\log\pi(X_{t})+\sqrt{2\rho(X_{t})}\circ\dot{W}$
in Stratonovitch form. The drift term in the Itô formulation (24) can be expressed as
(25) $\mathcal{L}\,\mbox{Id}=\rho\nabla\log\pi+\nabla\rho$
and, hence, the score function (12) is a finite $M$ approximation to
@@
However, recall that the recursive Langevin sampler (13) relies on a direct time-stepping approach to (24) and is based on
$m_{\epsilon}(x)=\mathcal{X}p_{\epsilon}(x)\approx\exp(\epsilon\mathcal{L})\,\mbox{Id}(x)$
instead of an approximation of the drift term on the right hand side of (24) via (25) followed by Euler–Maruyama. However, both approaches are formally consistent in the limit $\epsilon\to 0$ as
@@
While we have focused on the particular choice $K(x)=\rho(x)I$ in this section for simplicity, the discussion naturally extends to symmetric positive definite matrices $K(x)$.
5. Conditional sampling
In this section, we explore an extension of the sampling scheme (3.2) to the context of conditional generative modeling. More specifically, consider a random variable in $x=(y,z)$, which we condition on the second component for given $z=z^{\ast}$. In other words, we wish to sample from $\pi(y|z^{\ast})$ given samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from the joint distribution $\pi(x)=\pi(y,z)$.
Bayesian inference arises as a popular application of conditional sampling. In the traditional Bayesian inference framework, to generate samples from the posterior distribution, one typically requires access to both the prior and the likelihood, namely,
$\hat{\pi}(y):=\pi(y|z^{\ast})\propto\pi(z^{\ast}|y)\,\pi(y).$
However, conditional sampling provides a way to sample from the posterior just via the samples from the joint, obviating the need for the Bayesian update. In order to perform the required conditional sampling, we propose a method which combines approximate Bayesian computation (ABC) with our diffusion map based sampling algorithm.
Let us assume that we can generate $M$ samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from a joint distribution $\pi(y,z)$, which we then wish to condition on a fixed $z^{\ast}$. As before, we construct vectors of transition probabilities $p_{\epsilon}(x)\in\mathbb{R}^{M}$ based on the samples $\{x^{(i)}\}_{i=1}^{M}$. We assume that the bandwidth parameter $\epsilon$ used in the diffusion map approximation is also applied in the ABC misfit function, that is,

### v2

@@ 3. Approximating the conditional mean
(10) $p_{\epsilon}(x)=\frac{D(v_{\epsilon})t_{\epsilon}(x)}{v_{\epsilon}^{\top}t_{\epsilon}(x)}\in\mathbb{R}^{M}$
for all $x\in\mathbb{R}^{d}$. This vector gives the transition probabilities from $x$ to $\{x^{(i)}\}_{i=1}^{M}$ and provides a finite-dimensional approximation to the conditional probability distribution $\pi_{\epsilon}(\cdot|x)$ of the true underlying diffusion process; i.e., the semigroup $\exp(\epsilon\mathcal{L})$ with generator $\mathcal{L}$ corresponding to the generalized reversible diffusion process (6). See Section 4 for more details.
Finally, using the probability vector $p_{\epsilon}$ and introducing the data matrix of samples
(11) $\mathcal{X}=(x^{(1)},\ldots,x^{(M)})\in\mathbb{R}^{d\times M},$
our sample-based approximation of the conditional mean is provided by
@@
3.1. Sampling algorithms
We now present our main Langevin sampling strategies based on the previously introduced probability vector $p_{\epsilon}(x)$ in (10). Sampling schemes of the form (4) have the same drift term $m_{\epsilon}(x)$, but differ in the way the diffusion matrix $\Sigma(x)$ is defined. We consider a data-unaware diffusion as well as a data-aware diffusion which turns out to be advantageous in generating new samples from the data distribution $\pi$. See the numerical experiments in Section 6.
3.1.1. Langevin sampler with data-unaware diffusion
@@
(13) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m_{\epsilon}(X_{n})-X_{n}}{\epsilon}\right)+\sqrt{2K(X_{n})}\Xi_{n}$
as an approximation to (6), where $\Delta\tau$ is the time step and $\Xi_{n}\sim{\rm N}(0,\Delta\tau I)$. If $K=I$, we obtain the score function approximation
@@
in (1). Furthermore, by taking $\Delta\tau=\epsilon$, we have
(15) $X_{n+1}=m_{\epsilon}(X_{n})+\sqrt{2K(X_{n})}\Xi_{n}.$
Note that (15) fits into the general formulation (4) with $\Sigma(x)=2K(x)$.
Let us briefly discuss the qualitative behavior of the time-stepping method (15) as a function of $\epsilon>0$. For large $\epsilon$, the expected value $m_{\epsilon}(x)$ will become essentially independent of the current state $X_{n}$ and the diffusion process will sample from a centred Gaussian. For $\epsilon\to 0$, on the other hand, the probability vector $p_{\epsilon}(x)$ can potentially degenerate into a vector with a single entry approaching one with all other entries essentially becoming zero. Hence a key algorithmic challenge is to find a good value for $\epsilon$ and a suitable $K(x)$, which guarantee both good mixing and accuracy, i.e., $X_{n}\sim\pi$ as $n\to\infty$.
@@ 3.1.2. Langevin sampler with data-aware diffusion
From (10) and (11), one can also define the scaled conditional covariance matrix,
(16) $C(x)=\epsilon^{-1}(\mathcal{X}-m_{\epsilon}(x)1_{M}^{\top})D(p_{\epsilon}(x))(\mathcal{X}-m_{\epsilon}(x)1_{M}^{\top})^{\top}\in\mathbb{R}^{d\times d},$
which is the (scaled) covariance matrix associated with the probability vector $p_{\epsilon}(x)$. Here $1_{M}\in\mathbb{R}^{M}$ denotes the $M$-dimensional vector of ones. Therefore, one can more directly implement a Gaussian approximation associated with the transition probabilities $p_{\epsilon}(x)$ and introduce the update
(17) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m_{\epsilon}(X_{n})-X_{n}}{\epsilon}\right)+\sqrt{C(X_{n})}\Xi_{n}.$
@@
(18) $X_{n+1}=m_{\epsilon}(X_{n})+\sqrt{C(X_{n})}\Xi_{n},$
which we found to work rather well in our numerical experiments since it directly captures the uncertainty contained in the data-driven coupling $P_{\epsilon}$. The scheme (18) corresponds to setting $\Sigma(x)=C(x)$ in (4). Also note that the scheme (18) still depends on $K(x)$ through the probability vector $p_{\epsilon}(x)$.
@@ 3.2. Algorithmic properties
The Lemma follows from the definition (12), which we write as $m_{\epsilon}(x)=\sum_{j=1}^{M}x^{(j)}p_{\epsilon_{j}}(x)$, and the fact that $p_{\epsilon}(x)$ is a probability vector with $0\leq p_{\epsilon_{j}}(x)\leq 1$ for all $\epsilon>0$ and all $x\in\mathbb{R}^{d}$. ∎
This establishes stability of the Langevin samplers (15) and (18) for all step-sizes $\epsilon>0$. The next lemma shows that the Langevin sampler (15) is geometrically ergodic.
@@
$\mathcal{C}=\{x:\|x\|\leq R\}$
for suitable $R>0$. Since $m_{\epsilon}(X_{n})\in\mathcal{C}_{M}$ and $\pi$ has compact support, one can find a radius $R>0$, which is independent of the training data $\{x^{(i)}\}_{i=1}^{M}$, such that $\mathcal{C}_{M}\subset\mathcal{C}$ and
@@
for all $X_{n}\notin\mathcal{C}$ with $10\leq\lambda<$ 1. Furthermore, because of the additive Gaussian noise in (15), there is a probability density function $\nu(x)$ and a constant $\delta>0$ such that
@@
for all $x,x^{\prime}\in\mathcal{C}$. Here ${\rm n}(x;m,\Sigma)$ denotes the Gaussian probability density function with mean $m$ and covariance matrix $\Sigma$. In other words, $\mathcal{C}$ is a small set in the sense of [38]. Geometric ergodicity follows from Theorem 15.0.1 in [38]. See also the self-contained presentation in [36]. ∎
We note that extending Lemma 3.2 to the time-stepping scheme (18) with a data-aware diffusion is non-trivial since the covariance matrix (16) may become singular.
Lemma 3.1 suggests to replace the sampling step (15) by the associated split-step scheme
(20a) $X_{n+1/2}$ $=X_{n}+\sqrt{2K(X_{n})}\Xi_{n},$
(20b) $X_{n+1}$ $=\mathcal{X}p_{\epsilon}(X_{n+1/2}).$
@@ 3.3. Variable bandwidth diffusion
It is well-known from the literature on diffusion maps that a variable bandwidth can improve the approximation quality for fixed sample size $M$ [1]. Here we utilize the same idea. However, we no longer insist on approximating the standard generator with $K=I$, since we only wish to sample from the distribution $\pi$ rapidly. Hence, we consider reversible diffusion processes (6) with
@@
It is an active area of research to select a $\rho$ that increases the spectral gap of $\mathcal{L}$ while not increasing computational complexity. A larger spectral gap implies a faster convergence rate [43], indicating that the generated samples are closer to the reference at a finite time, exhibiting a high accuracy. We demonstrate numerically in Section 6 that $\rho$ can indeed be used to increase the sampling accuracy. More specifically, the bandwidth $\rho(x)$ is chosen as
@@
where $\beta\leq 0$ is a parameter and the unknown sampling distribution $\pi$ is approximated by an inexpensive low accuracy density estimator [46, 52]. One finds that the variable bandwidth parameter $\beta$ in (23) and the scaling parameter $\epsilon$ both influence the effective step-size in the Markov chain approximation (8) for (22). In order to disentangle the two scaling effects we modify the construction of the entries (7) in $T_{\epsilon}$ as follows. We first compute $\pi_{i}\approx\pi(x^{(i)})$ over all data points and then rescale these typically unnormalized densities:
@@
for $i=1,\ldots,M$, i.e., $K(x^{(i)})=\rho_{i}I$ in (7) and, more generally,
@@ 4. Connections to the generator of Langevin dynamics
In this section, we discuss the proposed methodology in the limit $M\to\infty$ under the idealised assumption that the target distribution $\pi$ is in fact known. It is well-known that $K=I$ implies that the Schrödinger bridge approximates the semi-group corresponding to the generator of standard Langevin dynamics [34, 54]. Employing a variable bandwidth (22) instead alters the semi-group and thus the underlying generator. However, since in both cases the density $\pi(x)$ remains invariant, this is not an issue as we are only concerned with drawing samples from the distribution and not in reproducing any dynamical features.
More precisely, using (22) in (7), we consider the associated Markov chain approximation given by (8). We formally take the limit $M\to\infty$ and denote the limiting operator by $\mathcal{P}_{\epsilon}$ [54]. One formally obtains
@@
(25) $\mathcal{L}f=\pi^{-1}\nabla\cdot(\pi\rho\nabla f)=\nabla\cdot(\rho\nabla f)+\rho\nabla\log\pi\cdot\nabla f.$
@@
$\dot{X}_{\tau}=\rho(X_{\tau})\left(\nabla(\log\pi(X_{\tau})-\log\mu(X_{\tau})\right),\qquad X_{0}\sim\mu_{0},$
@@
Please also note that the generator (25) corresponds to the diffusion process (6) with $K(x)=\rho(x)I$ in Itô form and to
$\dot{X}_{\tau}=\rho(X_{\tau})\nabla\log\pi(X_{\tau})+\sqrt{2\rho(X_{\tau})}\diamond\dot{W}_{\tau}$
in Klimontovitch form [23]. The drift term in the Itô formulation (6) can be expressed as
(26) $\mathcal{L}\,\mbox{Id}(x)=\rho\nabla\log\pi+\nabla\rho$
and, hence, (14) provides a finite $M$ approximation to the score function
@@
However, recall that the recursive Langevin sampler (15) relies on a direct time-stepping approach to (6) for $K(x)=\rho(x)I$ and is based on
$m_{\epsilon}(x)=\mathcal{X}p_{\epsilon}(x)\approx\exp(\epsilon\mathcal{L})\,\mbox{Id}(x),$
instead of an approximation of the drift term on the right hand side of (6) via (26) followed by a Euler–Maruyama time discretization. However, both approaches are formally consistent in the limit $\epsilon\to 0$ as
@@
While we have focused on the particular choice $K(x)=\rho(x)I$ in this section for simplicity, the discussion naturally extends to general symmetric positive definite matrices $K(x)$.
5. Bayesian inference and conditional sampling
We note that the previously discussed sampling methods can easily be extended to reversible diffusion processes of the form
(27) $\dot{X}_{\tau}=-\nabla V(X_{\tau})+\nabla\log\pi(X_{\tau})+\sqrt{2}\dot{W}_{\tau},$
where $V(x)$ denotes a known potential such as the negative log-likelihood function in case of Bayesian inference with $\pi$ taking the role of the prior distribution. We obtain, for example, the adjusted time-stepping scheme
$X_{n+1/2}$ $=X_{n}-\epsilon\nabla V(X_{n})+\sqrt{2}\,\Xi_{n},$
$X_{n+1}$ $=\mathcal{X}p_{\epsilon}(X_{n+1/2})$
for given samples $\{x^{(i)}\}_{i=1}^{M}$ from an unknown (prior) distribution $\pi(x)$. This numerical scheme approximately samples from the invariant distribution
$\tilde{\pi}(x)\propto e^{-V(x)}\pi(x).$
Remark 5.1.
We briefly divert to optimization of a regularised minimization problem of the form
$x^{\ast}=\arg\min_{x}\left\{V(x)-\log\pi(x)\right\}$
for given potential $V(x)$ and regulariser $-\log\pi(x)$. Again assuming that only samples $\{x^{(i)}\}_{i=1}^{M}$ of $\pi$ are available, an algorithm for approximating $x^{\ast}$ can be defined as follows:
$x_{n+1/2}$ $=x_{n}-\epsilon\nabla V(x_{n}),$
$x_{n+1}$ $=\mathcal{X}p_{\epsilon}(x_{n+1/2}).$
Here $m_{\epsilon}(x)=\mathcal{X}p_{\epsilon}(x)$ approximates the optimization update associated with the regulariser $-\log\pi(x)$. The iteration is stable and any limiting point $x_{\infty}$ is contained in the convex hull of the data points $\{x^{(i)}\}_{i=1}^{M}$.
Alternatively, the sampling schemes (20) and (21) can also easily be extended to conditional generative modeling. More specifically, consider a random variable in $x=(y,z)$, which we condition on the second component for given $z=z^{\ast}$. In other words, we wish to sample from $\pi(y|z^{\ast})$ given samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from the joint distribution $\pi(x)=\pi(y,z)$.
In order to perform the required conditional sampling, we propose a method which combines approximate Bayesian computation (ABC) with our diffusion map based sampling algorithm. Let us assume that we can generate $M$ samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from a joint distribution $\pi(y,z)$, which we then wish to condition on a fixed $z^{\ast}$. As before, we construct vectors of transition probabilities $p_{\epsilon}(x)\in\mathbb{R}^{M}$ based on the samples $\{x^{(i)}\}_{i=1}^{M}$. We assume that the bandwidth parameter $\epsilon$ used in the diffusion map approximation is also applied in the ABC misfit function, that is,